OpenStudy (anonymous):

BC trucking company charges $56.00(p) to deliver 8 cartons (c) to a nearby town and $78.00 to deliver 24 cartons find a linear equation usin (c,f) for these data points use these equation to find how much BC trucking company charges to deliver 72 cartons to the town

OpenStudy (amistre64):

(carton,fees) 24, 78 -8,-56 ------ 16,22 slope = y/x = 22/16 = 11/8

OpenStudy (amistre64):

f = (11/8)c + b 56 = 11(8)/8 + b 56 = 11 + b 45 = b the equation should be...barring any stupidity on my part... f = (11/8)c + 45

OpenStudy (amistre64):

when c = 72; f = 11(72)/8 + 45 f = 792/8 +45 f = 99 + 45 = $144

OpenStudy (anonymous):

thank you, very very very much!!!!